OpenStudy (sleepyjess):

\(\sf cot(x\dfrac{\pi}{2})=-tan~x\)

OpenStudy (turingtest):

what are we doin' here?

OpenStudy (sleepyjess):

Whoops, proving it \(\sf cot(x\dfrac{\pi}{2})=-tan~x\)

ganeshie8 (ganeshie8):

Looks there is a negative or positive sign missing on the left hand side inside cot

OpenStudy (sleepyjess):

Gah, it should be \(\sf cot(x-\dfrac{\pi}{2})=-tan~x\) sorry :/

OpenStudy (anonymous):

\(\text{co}\) is actually defined such that \(\text{co}f(x) = f(\pi/2-x)\)

OpenStudy (anonymous):

So we can start with: \[ \cot(x-\pi/2) = \tan(\pi/2-(x-\pi/2)) \]

OpenStudy (sleepyjess):

Ok, that makes sense

OpenStudy (anonymous):

I believe that simplifies to \(\tan(\pi-x)\).

OpenStudy (sleepyjess):

One question, why is it tan instead of -tan?

OpenStudy (anonymous):

The next step is probably the angle subtraction property.

OpenStudy (anonymous):

angle difference property of tangent.

OpenStudy (anonymous):

Hmm, but \(\tan(\pi)\) is undefined unfortunately...

OpenStudy (anonymous):

Wait, no, \(\tan\pi = 0\).

OpenStudy (anonymous):

\[ \tan(\pi-x)=\frac{\tan(\pi)-\tan(x)}{1+\tan(\pi)\tan(x)} = \frac{0-\tan(x)}{1+0} = -\tan(x) \]